OpenStudy (anonymous):

One of your friends sends you an email asking you to explain how all of the following expressions have the same answer. the cube root of x cubed :x to the one–third power • :x to the one–third power • x to the one–third power 1 over x to the –1 power the eleventh root of the quantity of x to the fifth times x to the fourth times x squared

OpenStudy (anonymous):

Can anyone help please ? i really suck at this

OpenStudy (triciaal):

rules of exponents labeling each section above as a, b, c d a) cubed root = raised to the power 1/3 (x^3)(1/3) = x^1 = x b) raise the power to a power multiply x^(1/3)^3= x^(1/3)(3) = x(1) = x c) 1/(x^-1 ) 1/ (1/x ) to divide by fraction invert and multiply (1)(x/1) = x d) (x^5)(x^4)(x^2)^(1/11) multiply with the same base add the exponents base = x 5 + 4 + 2 = x^11^1/11 = x ^1 = x

OpenStudy (anonymous):

thank you so much ! well explained. your awesome.

OpenStudy (triciaal):

your welcome